# Forte : Finding Outliers with Representation Typicality Estimation

### v1

@@ 1 Introduction
In this paper, we present the following key contributions:
@@
    Forte, a novel framework combining diverse representation learning techniques (CLIP, ViT-MSN, and DINOv2) for typicality estimation with non-parametric density estimator models (OCSVM, KDE, or GMM) to detect atypical samples, i.e., out-of-distribution (OOD) and synthetic images generated by foundation models. Forte requires no class labels, exposure to OOD data during training, or restrictions on the architecture of generative models.
@@ 2 Related Works
Figure 1: Visualization of Precision, Recall, Density, and Coverage metrics for reference and test samples in a 2D feature space, with nearest neighbour k=1. Precision is depicted by blue circles around reference samples, recall by red circles around test samples, density by teal circles around test samples, and coverage by green circles around test samples.
3 Methodology
@@ 3.1 Problem Setup
We start with data ${X_{i}=x}_{i}^{n}$ that comes from a true but unknown distribution $p$, where each $X_{i}$ is an independent and identically distributed random variable from $p$ and each $x_{i}$ is an observation. In the context of out-of-distribution detection, we consider that the unseen data might come from a mix of the true distribution $p$ and an unknown confounding distribution $\tilde{p}$ (which can be from an OOD benchmark or synthetic data generated using a model such as Stable Diffusion). The mixed distribution is denoted by $\grave{X}\sim\alpha p(\grave{X})+(1-\alpha)\tilde{p}(\grave{X})$, where $\alpha$ is an unknown mixing parameter. Since both $\alpha$ and $\tilde{p}$ are unknown, we cannot directly obtain samples from $\tilde{p}$ or make any assumptions about $\alpha$ and $\tilde{p}$. The goal of OOD detection is to develop a decision rule to determine when the input data $\grave{X}$ is atypical.
@@ 3.2 Creating Generalized Novel Summary Statistics
We propose the following per-point metrics, based on previous work in manifold estimation inside representation spaces by Naeem et al. (2020b) and Kynkäänniemi et al. (2019a) for capturing different facets of the generated samples. Let $\mathds{1}(\cdot)$ be the indicator function, $S({x_{j}^{r}}{j=1}^{m})=\bigcup{j=1}^{m}B(x_{j}^{r},\mathrm{NND}_{k}(x_{j}^{r}))$, where $B(x,r)$ is a Euclidean ball centered at $x$ with radius $r$, and $\mathrm{NND}k(x_{j}^{r})$ is the distance between $x_{j}^{r}$ and its $k$-th nearest neighbor in ${x_{j}^{r}}{j=1}^{m}$, excluding itself.
@@
    $\mathrm{precision_{pp}^{(i)}}=\mathds{1}\left(x_{i}^{g}\in S({x_{j}^{r}}{j=1}^{m})\right).$ (1)
@@
    $\mathrm{recall_{pp}^{(i)}}=\frac{1}{m}\sum{j=1}^{m}\mathds{1}\left(x_{j}^{r}\in B(x_{i}^{g},\mathrm{NND}_{k}(x_{i}^{g}))\right).$ (2)
@@
    Density per point is computed for each test point by counting the number of reference points within its nearest neighbor distance and dividing by the product of $\mathrm{nearest}_{k}$ and the total number of reference points:
    $\mathrm{density_{pp}^{(i)}}=\frac{1}{km}\sum{j=1}^{m}\mathds{1}\left(x_{j}^{r}\in B(x_{i}^{g},\mathrm{NND}_{k}(x_{i}^{g}))\right).$ (3)
@@
    $\mathrm{coverage_{pp}^{(i)}}=\mathds{1}\left(\min{j=1,\dots,m}d(x_{i}^{g},x_{j}^{r})<\mathrm{NND}_{k}(x_{i}^{g})\right).$ (4)
@@
We extract representations from reference and test images using CLIP Radford et al. (2021b), which learns to map images and text into a shared latent space, self-supervised models Gui et al. (2023) like ViT-MSN Assran et al. (2022), which predicts similarity between masked views of the same image, and DiNo v2 Oquab et al. (2023a), which distills knowledge from a teacher to a student network. Combining representations from these diverse models captures distinct aspects of the data via different manifolds and improves robustness of our anomaly detection approach, as we observe later in Table 3.
@@ 4 Results
Table 1 & 2 present the performance comparison between our proposed methods (Forte+SVM, Forte+KDE, and Forte+GMM) and various state of the art supervised and unsupervised techniques for out-of-distribution (OOD) detection techniques. Our methods consistently outperform techniques across all challenging dataset pairings, demonstrating their effectiveness in detecting OOD samples without relying on labeled data.
@@
Table 1: Performance comparison between our method and established supervised SoTA methods on various out-of-distribution (OOD) detection tasks. Our method consistently achieves the best performance across a range of dataset pairings, particularly outperforming on challenging datasets like NINCO and SSB-Hard. It also excels at detecting covariate shifted ID datasets. Key configurations include ViT-B trained with Cross Entropy and the RMDS postprocessor, and DINOv2+MLS, which utilizes training with Linear Probe DINOv2 alongside the MLS postprocessor. For reliable measurements, Forte is run with 10 random seeds.
Table 2: Comparison of performance figures between our method and various unsupervised baselines for out-of-distribution (OOD) detection tasks, demonstrating superior performance of our method across all challenging dataset pairings. For reliable measurements, Forte is run with 10 random seeds.
@@ 4.1 Performance on Synthetic Data Detection
To generate synthetic data for assessing distributional robustness, we first employ the Stable Diffusion Img2Img setting (Rombach et al., 2022), where a diffusion-based model can generate new images conditioned on an input image and a text prompt. We use the Stable Diffusion 2.0 base model and generate images with varying strength parameters (0.3, 0.5, 0.7, 0.9, 1.0) to control the influence of the input image on the generated output, essentially allowing our real distribution to be prior of controllable strength for the generated distribution. We also use the Stable Diffusion 2.0 text-to-image model to generate new images conditioned on the captions generated by BLIP(Li et al., 2022) for each real image from the reference distribution. This allows us to create images that are semantically similar to the real images but with novel compositions and variations. Finally, we also generate images directly from the class name associated with each real image (e.g., "a photo of a {monarch butterfly}, in a natural setting"). This provides a baseline for generating images that capture the essential characteristics of the class without relying on specific input images or captions. The image generation pipeline is implemented using the Hugging Face Transformers (Wolf et al., 2020) and Diffusers (von Platen et al., 2022) libraries, which provide high-level APIs for working with pre-trained models. Examples can be found in Figure 2 Figure 3 Figure 4
@@ 4.2 Performance on Medical Image Datasets
Table 5: Out-of-distribution (OOD) detection using Forte, applied to medical image datasets. Strong performance by Forte suggests the presence of batch effects and a need for data harmonization. Refer to Appendix D for more details on the FastMRI and OAI datasets. For reliable estimation, performance is measured over 10 random seeds.
@@ Appendix C Appendix : Theoretical Properties of per-point PRDC
    $\text{R}(\mathbf{x}^{\prime})=\frac{1}{N_{\text{train}}}\sum_{\mathbf{x}\in X_{\text{ID}}^{\text{train}}}\mathbb{I}\left[d(\mathbf{x}^{\prime},\mathbf{x})\leq r_{k}(\mathbf{x})\right].$
@@
    $\text{R}(\mathbf{x}^{\prime})=\frac{1}{N_{\text{train}}}\sum_{\mathbf{x}\in X_{\text{ID}}^{\text{train}}}\mathbb{I}\left[d(\mathbf{x}^{\prime},\mathbf{x})\leq r_{k}(\mathbf{x})\right].$
  • From our earlier results for precision, we have that the expected value of the indicator function is $\frac{k}{N_{\text{train}}+1}$, yielding the Expected Value for Recall per point as:
    $\mathbb{E}[\text{R}(\mathbf{x}^{\prime})]=\frac{1}{N_{\text{train}}}\times N_{\text{train}}\times\frac{k}{N_{\text{train}}+1}=\frac{k}{N_{\text{train}}+1}.$
@@
  • Definition is similar to Precision per point. Therefore, Expected Coverage per point:
    $\mathbb{E}[\text{C}(\mathbf{x}^{\prime})]\approx 1-e^{-k}.$
@@
    Similarly for Coverage per point.
@@
    $p=\frac{k}{N_{\text{train}}+1}$
@@ C.3 Key Properties and Characteristics of the Forte Algorithm
Appendix D Appendix: Medical Image Datasets

### v2

@@ 1 Introduction
In this paper, we hypothesize that many of the shortcomings with typicality-based approaches could be addressed using statistics which tune to the semantic content of the data. We propose to leverage self-supervised representations, which extract semantic information while discarding many potential confounding features (e.g. textures, backgrounds). Our specific contributions are:
@@
    Forte, a novel framework combining diverse representation learning techniques (CLIP, ViT-MSN, and DINOv2) for typicality estimation, uses both parametric (GMM) and non-parametric (KDE, OCSVM) density-based methods to detect atypical samples, i.e., out-of-distribution (OOD) and synthetic images generated by foundation models. Forte requires no class labels, exposure to OOD data during training, or restrictions on the architecture of generative models.
@@ 2 Related Works
Figure 1: Visualization of Precision, Recall, Density, and Coverage metrics for reference and test samples in a 2D feature space, with nearest neighbour k=1. Precision is depicted by blue circles around reference samples, recall by red circles around test samples, density by solid teal circles around reference samples, and coverage by green circles around test samples.
3 Forte: A Framework for OOD Detection Using Per-Point Metrics
In this section, we introduce Forte, a novel framework that combines diverse representation learning techniques with per-point summary statistics and non-parametric density estimation models to detect out-of-distribution (OOD) and synthetic data.
@@ 3.1 Problem Setup
We start with a dataset $X=\{x_{i}^{r}\}_{i=1}^{m}$ that comes from a true but unknown distribution $p$, where each $x_{i}\in\mathbb{R}^{d}$ is sampled independently and identically from $p$. In the context of out-of-distribution detection, we consider that the unseen data $\{x_{j}^{g}\}_{j=1}^{n}$ might come from a mix of the true distribution $p$ and an unknown confounding distribution $\tilde{p}$ (which can be from an OOD benchmark or synthetic data generated using a model such as Stable Diffusion). The mixed distribution is denoted by $\grave{X}\sim\alpha p(\grave{X})+(1-\alpha)\tilde{p}(\grave{X})$, where $\alpha$ is an unknown mixing parameter. Since both $\alpha$ and $\tilde{p}$ are unknown, we cannot directly obtain samples from $\tilde{p}$ or make any assumptions about $\alpha$ and $\tilde{p}$. The goal of OOD detection is to develop a decision rule to determine when the input data $\grave{X}$ is atypical.
@@ 3.2 Creating Generalized Novel Summary Statistics
We propose the following per-point metrics, novelly adopted to OOD detection from previous work in manifold estimation inside representation spaces by Naeem et al. (2020b) and Kynkäänniemi et al. (2019a) for capturing different facets of the generated samples. Let $\mathds{1}(\cdot)$ be the indicator function, $S(\{x_{i}^{r}\}_{i=1}^{m})=\bigcup_{i=1}^{m}B(x_{i}^{r},\mathrm{NND}_{k}(x_{i}^{r}))$, where $B(x,r)$ is a Euclidean ball centered at $x$ with radius $r$, and $\mathrm{NND}_{k}(x_{i}^{r})$ is the distance between $x_{i}^{r}$ and its $k$-th nearest neighbor in $\{x_{i}^{r}\}_{i=1}^{m}$, excluding itself.
@@
    $\mathrm{precision_{pp}^{(j)}}=\mathds{1}\left(x_{j}^{g}\in S(\{x_{i}^{r}\}_{i=1}^{m})\right).$ (1)
@@
    $\mathrm{recall_{pp}^{(j)}}=\frac{1}{m}\sum_{i=1}^{m}\mathds{1}\left(x_{i}^{r}\in B(x_{j}^{g},\mathrm{NND}_{k}(x_{j}^{g}))\right).$ (2)
@@
    Density per point is computed for each test point by counting the number of reference points for which it is within that points nearest neighbor distance and dividing by the product of $k$ and the total number of reference points:
    $\mathrm{density_{pp}^{(j)}}=\frac{1}{km}\sum_{i=1}^{m}\mathds{1}\left(x_{j}^{g}\in B(x_{i}^{r},\mathrm{NND}_{k}(x_{i}^{r}))\right).$ (3)
@@
    $\mathrm{coverage_{pp}^{(j)}}=\mathds{1}\left(\min_{i}(d(x_{j}^{g},x_{i}^{r}))<\mathrm{NND}_{k}(x_{j}^{g})\right).$ (4)
@@
These per-point metrics serve as summary statistics that capture local geometric properties of the data manifold in the feature space. They enable us to model the distribution of in-distribution (ID) data and identify out-of-distribution (OOD) samples effectively. We extract representations from reference and test images using CLIP Radford et al. (2021b), which learns to map images and text into a shared latent space, self-supervised models Gui et al. (2023) like ViT-MSN Assran et al. (2022), which predicts similarity between masked views of the same image, and DiNo v2 Oquab et al. (2023a), which distills knowledge from a teacher to a student network. Combining representations from these diverse models captures distinct aspects of the data via different manifolds and improves robustness of our anomaly detection approach, as we observe later in Table 3.
@@ 5 Results
Tables 1 & 2 present the performance comparison between our proposed methods (Forte+SVM, Forte+KDE, and Forte+GMM) and various state of the art supervised and unsupervised techniques for out-of-distribution (OOD) detection techniques. Our methods consistently outperform techniques across all challenging dataset pairings, demonstrating their effectiveness in detecting OOD samples without relying on labeled data.
@@
Table 1: Performance comparison between our method and established supervised SoTA methods on various out-of-distribution (OOD) detection tasks. Our method consistently achieves the best performance across a range of dataset pairings, particularly outperforming on challenging datasets like NINCO and SSB-Hard. It also excels at detecting covariate shifted ID datasets. Key configurations include ViT-B trained with Cross Entropy and the RMDS postprocessor, and DINOv2+MLS, which utilizes training with Linear Probe DINOv2 alongside the MLS postprocessor. For reliable measurements, Forte is run with 10 random seeds. More comparisons are provided in Fig. 16
Table 2: Comparison of performance figures between our method and various unsupervised baselines for out-of-distribution (OOD) detection tasks, demonstrating superior performance of our method across all challenging dataset pairings. For reliable measurements, Forte is run with 10 random seeds. This is presented as a visualization in Fig.15
@@ 5.1 Performance on Synthetic Data Detection
To generate synthetic data for assessing distributional robustness, we first employ the Stable Diffusion Img2Img setting (Rombach et al., 2022), where a diffusion-based model can generate new images conditioned on an input image and a text prompt. We use the Stable Diffusion 2.0 base model and generate images with varying strength parameters (0.3, 0.5, 0.7, 0.9, 1.0) to control the influence of the input image on the generated output, essentially allowing our real distribution to be prior of controllable strength for the generated distribution. We also use the Stable Diffusion 2.0 text-to-image model to generate new images conditioned on the captions generated by BLIP(Li et al., 2022) for each real image from the reference distribution. This allows us to create images that are semantically similar to the real images but with novel compositions and variations. Finally, we also generate images directly from the class name associated with each real image (e.g., "a photo of a {monarch butterfly}, in a natural setting"). This provides a baseline for generating images that capture the essential characteristics of the class without relying on specific input images or captions. The image generation pipeline is implemented using the Hugging Face Transformers (Wolf et al., 2020) and Diffusers (von Platen et al., 2022) libraries, which provide high-level APIs for working with pre-trained models. Examples can be found in Figure 2, Figure 3 & Figure 4.
@@ 5.2 Performance on Medical Image Datasets
Table 5: Out-of-distribution (OOD) detection using Forte, applied to medical image datasets. Strong performance by Forte suggests the presence of batch effects and a need for data harmonization. Refer to Appendix E for more details on the FastMRI and OAI datasets. For reliable estimation, performance is measured over 10 random seeds.
@@ 6 Discussion
DoSE (Morningstar et al., 2021) pioneered chaining multiple summary statistics for typicality measurement, using ID sample distributions to construct typicality estimators rather than direct statistic values. While groundbreaking, DoSE’s reliance on generative model likelihoods proved problematic, as subsequent work (Caterini & Loaiza-Ganem, 2022; Zhang et al., 2021) showed these can be unreliable for OOD detection. Our approach addresses these limitations through four key improvements: (1) utilizing self-supervised representations to capture semantic features, (2) incorporating manifold estimation to account for local topology, (3) unifying typicality scoring and downstream prediction models to minimize deployment overhead, and (4) eliminating additional model training requirements. These advances yield substantial empirical gains. While building upon DoSE’s fundamental statistical machinery, our modifications dramatically enhance practical performance.
@@ Appendix C Appendix : Theoretical Properties of per-point PRDC
    $\text{R}(\mathbf{x}^{\prime})=\frac{1}{N_{\text{train}}}\sum_{\mathbf{x}\in X_{\text{ID}}^{\text{train}}}\mathbb{I}\left[d(\mathbf{x}^{\prime},\mathbf{x})\leq r_{k}(\mathbf{x}^{\prime})\right].$
@@
    $\text{D}(\mathbf{x}^{\prime})=\frac{1}{kN_{\text{train}}}\sum_{\mathbf{x}\in X_{\text{ID}}^{\text{train}}}\mathbb{I}\left[d(\mathbf{x}^{\prime},\mathbf{x})\leq r_{k}(\mathbf{x})\right].$
@@
    $\text{R}(\mathbf{x}^{\prime})=\frac{1}{N_{\text{train}}}\sum_{\mathbf{x}\in X_{\text{ID}}^{\text{train}}}\mathbb{I}\left[d(\mathbf{x}^{\prime},\mathbf{x})\leq r_{k}(\mathbf{x}^{\prime})\right].$
  • Understanding $r_{k}(\mathbf{x}^{\prime})$: For each $\mathbf{x}^{\prime}$, $r_{k}(\mathbf{x}^{\prime})$ is the distance to its $k$-th nearest neighbor among the $N_{\text{train}}$ training samples. The distances $d_{i}=d(\mathbf{x}^{\prime},\mathbf{x}_{i})$ for $i=1,\dots,N_{\text{train}}$ are independent and identically distributed (i.i.d.) random variables because $\mathbf{x}^{\prime}$ and $\mathbf{x}_{i}$ are independent samples from the same distribution.
  • Probability that a Training Sample is within $r_{k}(\mathbf{x}^{\prime})$: The variable $\mathbb{I}\left[d(\mathbf{x}^{\prime},\mathbf{x}_{i})\leq r_{k}(\mathbf{x}^{\prime})\right]$ is an indicator that is 1 if $\mathbf{x}_{i}$ is among the $k$ nearest neighbors of $\mathbf{x}^{\prime}$ in the training set. The probability that a particular training sample $\mathbf{x}_{i}$ is within the $k$-nearest neighbors of $\mathbf{x}^{\prime}$ is:
    $P\left(d(\mathbf{x}^{\prime},\mathbf{x}_{i})\leq r_{k}(\mathbf{x}^{\prime})\right)=\frac{k}{N_{\text{train}}}$
    This is since the $k$ nearest neighbors are chosen among $N_{\text{train}}$ samples, each training sample has an equal chance of $\frac{k}{N_{\text{train}}}$ of being among the closest $k$ samples to $\mathbf{x}^{\prime}$.
    $\mathbb{E}\left[\mathbb{I}\left[d(\mathbf{x}^{\prime},\mathbf{x}_{i})\leq r_{k}(\mathbf{x}^{\prime})\right]\right]=\frac{k}{N_{\text{train}}}$
    $\mathbb{E}[\text{R}(\mathbf{x}^{\prime})]=\frac{1}{N_{\text{train}}}\sum_{i=1}^{N_{\text{train}}}\mathbb{E}\left[\mathbb{I}\left[d(\mathbf{x}^{\prime},\mathbf{x}_{i})\leq r_{k}(\mathbf{x}^{\prime})\right]\right]=\frac{1}{N_{\text{train}}}\times N_{\text{train}}\times\frac{k}{N_{\text{train}}}=\frac{k}{N_{\text{train}}}$
    $\mathbb{E}[\text{R}(\mathbf{x}^{\prime})]=\frac{k}{N_{\text{train}}}$
@@
  • Given both $\mathbf{x}^{\prime}$ and the training samples $\mathbf{x}$ are drawn from the same distribution $\mathcal{N}(\bm{\mu}_{\text{ID}},\sigma^{2}I)$.
  • The minimum distance $\min_{\mathbf{x}_{i}}d(\mathbf{x}^{\prime},\mathbf{x}_{i})$ is the smallest among $N_{\text{train}}$ distances drawn from the chi-squared distribution.
  • Distance to $k$-th Nearest Neighbor $r_{k}(\mathbf{x}^{\prime})$ is $r_{k}(\mathbf{x}^{\prime})$ is the $k$-th order statistic of the distances $d(\mathbf{x}^{\prime},\mathbf{x}_{i})$. We are interested in the probability:
    $P\left(\min_{\mathbf{x}_{i}}d(\mathbf{x}^{\prime},\mathbf{x}_{i})\leq r_{k}(\mathbf{x}^{\prime})\right)$
  • Since the minimum distance is always less than or equal to the $k$-th smallest distance, the event is always true:
    $\min_{\mathbf{x}_{i}}d(\mathbf{x}^{\prime},\mathbf{x}_{i})\leq r_{k}(\mathbf{x}^{\prime})$
    Therefore:
    $P\left(\min_{\mathbf{x}_{i}}d(\mathbf{x}^{\prime},\mathbf{x}_{i})\leq r_{k}(\mathbf{x}^{\prime})\right)=1$
    This is because the closest training sample to $\mathbf{x}^{\prime}$ is by definition among the $k$-nearest neighbors used to compute $r_{k}(\mathbf{x}^{\prime})$. Therefore:
    $\mathbb{E}[\text{C}(\mathbf{x}^{\prime})]=1$
@@
    The variance of coverage per point for ID samples is zero. This reflects the certainty that ID test samples will always satisfy the coverage condition.
@@
    $p=\frac{k}{N_{\text{train}}}$
@@ C.3 Key Properties and Characteristics of the Forte Algorithm
Appendix D Appendix : Ablation Study on Encoders
Insight into encoders: The results presented in Table. 3 demonstrate that richer representations significantly improve OOD detection performance, as evidenced by the ranking of encoders: CLIP > DINO v2 > MSN when used individually, and CLIP + DINO v2 > CLIP + MSN > DINO v2 + MSN in the two-model combinations. To further investigate this phenomenon, we conducted additional experiments. Specifically, we included DeIT with both ViT-B (Base) and ViT-Ti (Tiny) models and evaluated their OOD detection performance under the settings studied in Tables 2. These results, show that DeIT-B achieves 0.87 AUROC on CIFAR-100, while DeIT-Ti achieves 0.82 AUROC. This aligns with our hypothesis that more informative representations are essential for effective OOD detection. DeIT, trained with an objective approximating supervision, produces less informative embeddings compared to self-supervised encoders like DINO v2. Similarly, the DeIT-Ti model performs worse due to its reduced capacity for generating robust representations. We think these findings provide valuable insights into the utility of different encoders for OOD detection and offer guidance for practitioners seeking optimal performance.
Table 6: Comparison of AUROC and FPR95 performance figures for Base and Tiny DeIT models across the tasks in
Model In-Dist OOD Dataset AUROC FPR95
Base-DeIT CIFAR-10 CIFAR-100 0.8712 0.9926
Tiny-DeIT CIFAR-10 CIFAR-100 0.8261 0.9903
Base-DeIT CIFAR-10 SVHN 0.9554 0.4604
Tiny-DeIT CIFAR-10 SVHN 0.9296 0.6195
Base-DeIT CIFAR-10 Celeb-A 0.9871 0.0015
Tiny-DeIT CIFAR-10 Celeb-A 0.9929 0.0007
Appendix E Appendix: Medical Image Datasets
@@
Appendix F Pseudocode
Algorithm 1 OOD Detection Using Per-Point PRDC Metrics in Forte
1:Reference data features $\{x_{j}^{r}\}_{j=1}^{m}$
2:Test data features $\{x_{i}^{g}\}_{i=1}^{n}$
3:Number of nearest neighbors $k$
4:OOD detection performance metrics: AUROC, FPR@95
5:Feature Extraction (Preprocessing):
6: Use pre-trained models (e.g., CLIP, ViT-MSN, DINOv2) to extract features for both reference and test data.
7: Reference features: $\{x_{j}^{r}\}_{j=1}^{m}$
8: Test features: $\{x_{i}^{g}\}_{i=1}^{n}$
9:Compute Nearest Neighbor Distances:
10:for $j=1$ to $m$ do
11:     Compute $\mathrm{NND}_{k}(x_{j}^{r})$: distance to its $k$-th nearest neighbor in $\{x_{h}^{r}\}_{h=1,h\neq j}^{m}$.
12:end for
13:for $i=1$ to $n$ do
14:     Compute $\mathrm{NND}_{k}(x_{i}^{g})$: distance to its $k$-th nearest neighbor in $\{x_{h}^{g}\}_{h=1,h\neq i}^{n}$.
15:end for
16:Compute Per-Point PRDC Metrics for Test Data:
17:for $i=1$ to $n$ do
18:     Compute per-point metrics for $x_{i}^{g}$ relative to $\{x_{j}^{r}\}_{j=1}^{m}$:
  1. 1.
    Precision per point:
    $\mathrm{precision}_{pp}^{(i)}=\mathds{1}\left(x_{i}^{g}\in S(\{x_{j}^{r}\}_{j=1}^{m})\right)$
    where $S(\{x_{j}^{r}\}_{j=1}^{m})=\bigcup_{j=1}^{m}B\left(x_{j}^{r},\mathrm{NND}_{k}(x_{j}^{r})\right)$.
  2. 2.
    Recall per point:
    $\mathrm{recall}_{pp}^{(i)}=\frac{1}{m}\sum_{j=1}^{m}\mathds{1}\left(x_{j}^{r}\in B\left(x_{i}^{g},\mathrm{NND}_{k}(x_{i}^{g})\right)\right)$
  3. 3.
    Density per point:
    $\mathrm{density}_{pp}^{(i)}=\frac{1}{km}\sum_{j=1}^{m}\mathds{1}\left(x_{i}^{g}\in B\left(x_{j}^{r},\mathrm{NND}_{k}(x_{j}^{r})\right)\right)$
  4. 4.
    Coverage per point:
    $\mathrm{coverage}_{pp}^{(i)}=\mathds{1}\left(\min_{j=1,\dots,m}\|x_{i}^{g}-x_{j}^{r}\|<\mathrm{NND}_{k}(x_{i}^{g})\right)$
19:     Assemble feature vector $\phi^{(i)}=\left[\mathrm{precision}_{pp}^{(i)},\mathrm{recall}_{pp}^{(i)},\mathrm{density}_{pp}^{(i)},\mathrm{coverage}_{pp}^{(i)}\right]$
20:end for
21:Prepare Reference Training and Validation Sets:
22:Split $\{x_{j}^{r}\}_{j=1}^{m}$ into training set $\{x_{j}^{\text{train}}\}$ and validation set $\{x_{j}^{\text{valid}}\}$.
23:Compute Per-Point PRDC Metrics for Reference Training Data:
24:for each $x_{j}^{\text{train}}$ do
25:     Compute per-point metrics $\phi_{\text{ref}}^{(j)}$ following similar steps as above, relative to $\{x_{h}^{\text{train}}\}_{h=1}^{m_{\text{train}}}$
26:end for
27:Train Anomaly Detection Models:
28:Use the reference per-point metrics $\{\phi_{\text{ref}}^{(j)}\}$ to train unsupervised anomaly detection models: One-Class SVM (OCSVM), Kernel Density Estimation (KDE), & Gaussian Mixture Model (GMM).
29:Evaluate Models on Test Data:
30:for $i=1$ to $n$ do
31:     Compute anomaly scores $s^{(i)}$ for $\phi^{(i)}$ using the trained models. Use it to assign Ground Truth Labels, and calculate AUROC & FPR@95
32:end for
Notes:
  • $\mathds{1}(\cdot)$ is the indicator function, returning 1 if the condition is true, and 0 otherwise.
  • $B(x,r)$ denotes a ball (in Euclidean space) centered at $x$ with radius $r$.
  • $\mathrm{NND}_{k}(x)$ is the distance from point $x$ to its $k$-th nearest neighbor.
  • $S(\{x_{j}^{r}\}_{j=1}^{m})=\bigcup_{j=1}^{m}B\left(x_{j}^{r},\mathrm{NND}_{k}(x_{j}^{r})\right)$ represents the union of balls around each reference point $x_{j}^{r}$ with radius $\mathrm{NND}_{k}(x_{j}^{r})$.
Appendix G Comparative Evaluation
To contextualize the strong state-of-the-art performance achieved by Forte, we present the following infographics showing the evolution of methods and their performance on benchmark tasks.
Figure 15: AUROC performance of various unsupervised OOD detection methods over years for the CIFAR-10 (In-distribution) and the CIFAR-100(OOD), Celeb-A(OOD) and SVHN(OOD) dataset. The figure illustrates the progression of anomaly detection techniques, with methods represented using unique markers and colors.
Figure 16: Comparison of AUROC performance across various supervised out-of-distribution detection methods and datasets from the OpenOOD leaderboard. The figure presents results for five datasets (NINCO, SSB-Hard, iNaturalist, Textures, and OpenImage-O) with each subplot showcasing the progression of AUROC scores over the years. Each method is represented with a unique marker and color. Our method "Forte + GMM" is highlighted for its superior performance, demonstrating strong state-of-the-art results across all datasets. The x-axis represents the publication year, while the y-axis denotes the AUROC (%) scores.
Appendix H Miscellaneous Related Works